# Beyond Needle(s) in the Embodied Haystack: Environment, Architecture, and Training Considerations for Long Context Reasoning

### v1

@@ 5.2 Interactive Evaluation in $\infty$-Thor
We conduct an interactive online evaluation using the AI2THOR simulator to measure agent performance on our long-horizon test set. Our experiments analyze reward accumulation across various context extension methods, different scaling factors, and multiple fine-tuning context lengths: 32K, 64K, and 130K.
Training. We fine-tune the LLaVA-OneVision 7B model on our training set while freezing the vision encoder. We used 8 H100 GPUs with both tensor parallelism shoeybi2020megatronlmtrainingmultibillionparameter and pipeline parallelism huang2019gpipeefficienttraininggiant for the 32K context size, while Context Parallelism is employed for training on larger context sizes (64K and 130K). Additional training specifics are available in the Appendix.
Figure 5: Agent’s reward across different experimental configurations. We compare (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, (c) fine-tuning with different context lengths using Context Parallelism, and (d-e) combinations of scaling during both training and inference. (f) summarizes the most effective strategies, highlighting that exposure to longer contexts during training significantly improves performance. Non-planner models cannot generate valid actions after around 250 steps ($\approx$376K in context size).
Plan-level Evaluation. We evaluate agent performance based on reward accumulation given previous states and actions. Evaluation is performed at the plan-level, where a plan represents a short sequence of actions aimed at a specific intermediate goal. For example, a “Go to location” plan comprises actions like Move Ahead, Rotate Right/Left, and Look Up/Down. Each trajectory is thus composed of multiple sequential plans. At each step, agents are presented with a task goal alongside the previous state-action sequences, predicting subsequent actions and interacting continuously with the environment until the plan completion. Further evaluation details are provided in the Appendix.
Results and Discussion. Figure 5 illustrates the comparative results across six experimental configurations. The Planner trajectory serves as the performance upper-bound. We address the following key questions through our experimental results:
Q. Does combining context extension methods during both training and inference provide additional benefits? Figures 5(d) and 5(e) illustrate experiments with scaling applied at both training and evaluation stages. Results indicate that additional scaling at evaluation after fine-tuning with scaled RoPE provides no further performance improvement and may degrade performance at shorter context lengths ($\leq$300K tokens).
@@ 6 Conclusion
We presented $\infty$-Thor, a new framework for long-horizon embodied tasks designed to advance long-context understanding in embodied AI. Our framework enables scalable synthesis of long, complex trajectories paired with GT action sequences, and supports both offline training and online interaction with the environment. As part of this framework, we introduced a novel embodied QA benchmark, Needle(s) in the Embodied Haystack, that challenges agents to reason over sparse, temporally distant visual evidence embedded within extended trajectories. To equip models for this setting, we explored architectural adaptations including interleaved Goal–State–Action modeling, context extension techniques such as YaRN and LongRoPE, and efficient fine-tuning via Context Parallelism. Our experiments demonstrate that exposure to longer contexts during training significantly improves model performance, and the limitation of existing context extension techniques struggle with long-context reasoning. We hope our framework and benchmark encourage further research into models capable of robust long-horizon reasoning under realistic, interactive environments.
@@ A.1.3 Cross-validation with Multimodal LLMs
To ensure the answerability and clarity of the generated QA pairs, we perform cross-validation using four powerful multimodal LLMs: LLaVA-OneVision 7B [19], Qwen2.5-VL 7B [3], Deepseek-VL 7B [23], and Pixtral 12B [1]. Each model is prompted with the GT images corresponding to the annotated QA steps and asked to answer the associated questions. Given their strong performance on standard visual QA tasks, we use these models to assess whether a question can be correctly answered or not. We keep only the QA pairs that are correctly answered by at least one of the four models, and discard those that fail across all models. This helps improve dataset quality and filtering out ambiguous or visually ungroundable questions. Table 4 shows the accuracy of each model on the finalized QA set when evaluated with GT images. Notably, even with access to GT images, all models struggle with questions requiring reasoning over three or more evidence steps. To maintain the benchmark’s difficulty and support evaluation of more capable models in future, we manually inspect the multi-clue questions and include those that are answerable.
@@ A.2 Constructing Long-Horizon Trajectories
We use a classical task planner, which operates over PDDL-defined domains [29], to generate a low-level action sequence for the sampled goal, and simulate this plan in an interactive environment. If the rollout fails (e.g., due to collisions, object occlusions, or unreachable conditions), we discard the sequence and re-sample from the task pool. Otherwise, the successful rollout is retained and appended to the ongoing trajectory.
@@
1:Input: Task Pool $\mathcal{T}$, max sub goals $N$
@@ B.1 Interactive Evaluation in $\infty$-Thor
Appendix C Limitations
While $\infty$-Thor enables the generation of arbitrarily long trajectories, the diversity of environment layouts in AI2-THOR is inherently limited. This can lead to repetitive agent behaviors within certain scenes. For instance, compared to kitchen and living room scenes, bedroom scenes tend to involve fewer action types, mostly constrained to simple pick-and-place tasks within small spatial areas. Due to the limited scene size and low task diversity, we excluded bathroom scenes from our dev and test sets. In future work, we plan to integrate $\infty$-Thor with ProcTHOR [9], which supports procedurally generated environments, enabling a broader range of dynamic and diverse scene configurations.
@@ Appendix C Limitations
Finally, inference with LLMs in long-context settings remains a computational bottleneck, particularly as context lengths approach 1M tokens. Since our long-horizon tasks require access to information spread throughout the entire trajectory, full-context inference becomes increasingly expensive, even with Context Parallelism. One promising direction is to equip agents with memory systems that selectively keep relevant information from prior steps, allowing the model to reason without reprocessing the full context at each step. Other architectural solutions, such as sparse attention mechanisms [13, 16] or state-space models with linear-time inference [12], also hold potential for scalable long-context reasoning.

### v2

@@ 5.2 Interactive Evaluation in $\infty$-Thor
To measure agent performance on our long-horizon test set, we conduct an interactive online evaluation using the AI2THOR for high-level planning tasks and the ManipulaTHOR for low-level manipulation tasks.
High-level Planning Task. We fine-tune the LLaVA-OneVision 7B model on our training set while freezing the vision encoder. We used 8 H100 GPUs with both tensor parallelism (Shoeybi et al., 2020) and pipeline parallelism (Huang et al., 2019) for the 32K context size, while Context Parallelism is employed for training on larger context sizes (64K and 130K). Additional training specifics are available in the Appendix. Figure 5 presents the accumulated rewards over time across six experimental configurations. The Planner trajectory serves as the performance upper bound.
Figure 5: Agent’s reward across different experimental configurations for high-level planning tasks. We compare (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, and (c) fine-tuning with different context lengths using Context Parallelism. (d) summarizes the most effective strategies, highlighting that exposure to longer contexts during training significantly improves performance. Non-planner models cannot generate valid actions after around 250 steps ($\approx$376K in context size). More configurations are in Figure 7 in Appendix.
Figure 6: Agent’s reward with low-level manipulation tasks. Low-level VLA models are used to control the robot arm for Pick-up and Put actions.
In Figure 5(a), at a fixed scaling factor of x4, YaRN consistently achieves the highest performance, closely matching the Planner. Figure 5(b) shows that scaling YaRN further to x8 or x16 yields little additional gain. In Figure 5(c), fine-tuning with Context Parallelism extends context lengths to 64K and 130K tokens, enabling the model to learn much longer sequences (86 vs. 22 steps with 32K). Additional scaling at evaluation after fine-tuning (Figures 7(d,e) in Appendix) provides no improvement and may even degrade shorter-context performance. Overall, fine-tuning on long-trajectory datasets proves most effective (Figure 5(d)), while YaRN at x4 offers strong results when large-scale training data are unavailable.
Low-level Manipulation Task. Figure 6 shows agent rewards on low-level manipulation tasks using OpenVLA-7B (Kim et al., 2024) and SpatialVLA-4B (Qu et al., 2025). Ego-centric views are provided as image input, and the models generate the next action $(\Delta x,\Delta y,\Delta z)$ to control the robot arm. Both models perform below the Planner due to differences in robot arm configuration and out-of-distribution inputs, though SpatialVLA achieves slightly higher and more stable rewards than OpenVLA over long sequences. We provide more results and analysis in Appendix C.
@@ 6 Conclusion
We presented $\infty$-Thor, a new framework for long-horizon embodied tasks designed to advance long-context understanding in embodied AI. Our framework enables scalable synthesis of long, complex trajectories paired with high- and low-level action sequences, and supports both offline training and online interaction with the environment. As part of this framework, we introduced a novel embodied QA benchmark, Needle(s) in the Embodied Haystack, that challenges agents to reason over sparse, temporally distant visual evidence embedded within extended trajectories. To equip models for this setting, we explored architectural adaptations including interleaved Goal–State–Action modeling, context extension techniques such as YaRN and LongRoPE, and efficient fine-tuning via Context Parallelism. Our experiments demonstrate that exposure to longer contexts during training significantly improves model performance, and the limitation of existing context extension techniques struggle with long-context reasoning. We hope our framework and benchmark encourage further research into models capable of robust long-horizon reasoning under realistic, interactive environments.
Reproducibility statement
Experiments in this work were conducted using the publicly available simulators AI2-THOR and ManipulaTHOR, which allow for the deterministic replaying of all trajectories provided by $\infty$-Thor. We provide the complete source code, along with the generated datasets and configurations used for each experiment, and this will be made available in a public repository upon publication. Furthermore, all VLM models used in our experiments are open-source and available on Hugging Face. For the proprietary model Gemini Flash, we have declared the specific version of Gemini Flash used in the paper, and our experiments can be reproduced using the Google Cloud APIs. While the initial generation of trajectories involves some randomness in the sampling of target objects and destination positions, the resulting trajectories themselves are fully reproducible within the simulator environments. The detailed data processing steps and the experimental setup are further described in the Appendix.
@@ A.1.3 Cross-validation with Multimodal LLMs
To ensure the answerability and clarity of the generated QA pairs, we perform cross-validation using four powerful multimodal LLMs: LLaVA-OneVision 7B (Li et al., 2024a), Qwen2.5-VL 7B (Bai et al., 2025), Deepseek-VL 7B (Lu et al., 2024), and Pixtral 12B (Agrawal et al., 2024). Each model is prompted with the GT images corresponding to the annotated QA steps and asked to answer the associated questions. Given their strong performance on standard visual QA tasks, we use these models to assess whether a question can be correctly answered or not. We keep only the QA pairs that are correctly answered by at least one of the four models, and discard those that fail across all models. This helps improve dataset quality and filtering out ambiguous or visually ungroundable questions. Table 4 shows the accuracy of each model on the finalized QA set when evaluated with GT images. Notably, even with access to GT images, all models struggle with questions requiring reasoning over three or more evidence steps. To maintain the benchmark’s difficulty and support evaluation of more capable models in future, we manually inspect the multi-clue questions and include those that are answerable.
@@ A.2 Constructing Long-Horizon Trajectories
We use a classical task planner, which operates over PDDL-defined domains (Shridhar et al., 2020), to generate a low-level action sequence for the sampled goal, and simulate this plan in an interactive environment. If the rollout fails (e.g., due to collisions, object occlusions, or unreachable conditions), we discard the sequence and re-sample from the task pool. Otherwise, the successful rollout is retained and appended to the ongoing trajectory.
@@
1:Input: Task Pool $\mathcal{T}$, max sub goals NN
@@ B.1 Interactive Evaluation in $\infty$-Thor
Appendix C Additional Results
Figure 7: Agent’s reward across different experimental configurations: (a) context extension methods at fixed scaling (x4), (b) varying YaRN scaling factors, (c) fine-tuning with different context lengths using Context Parallelism, (d-e) combinations of scaling during both training and inference, and (f) summary of the most effective strategies.
C.1 Interactive Evaluation: High-level Planning
Results and Discussion. Figure 7 presents the accumulated rewards over time across six experimental configurations. The Planner trajectory represents the performance upper bound. Our analysis focuses on addressing the following key questions:
Q. Does combining context extension methods during both training and inference provide additional benefits? Results of experiments with scaling at both training and evaluation (Figures 7(d) and 7(e) in Appendix) indicate that additional scaling at evaluation after fine-tuning with scaled RoPE provides no further performance improvement and may degrade performance at shorter context lengths ($\leq$300K tokens).
C.2 Interactive Evaluation: Low-level Manipulation
Table 6: Success rates of OpenVLA-7B and SpatialVLA-4B on low-level Pick-up and Put tasks in ManipulaTHOR.
Task OpenVLA-7B SpatialVLA-4B
Pick-up 16.52% 18.70%
Put 2.08% 1.19%
We evaluate the ability of existing VLA models, OpenVLA-7B (Kim et al., 2024) and SpatialVLA-4B (Qu et al., 2025), to control robot arms on low-level manipulation tasks. The evaluation protocol follows the procedure described in Appendix B.1, with the only difference being that low-level VLA models are used to directly execute Pick-up and Put actions through arm control. For the text input, we provide task-specific instructions (“Pick-up” or “Put” object), since the existing models are only trained on single-task formulations. For the image input, we use an ego-centric camera view. Since these VLA models were originally trained on datasets with different viewing angles, there remains considerable room for improvement by incorporating additional camera perspectives during training and evaluation.
For the evaluation criterion, ManipulaTHOR implements a “magnet sphere” hand mechanism: if a pickupable object is within a specified radius of the agent’s hand when the PickupObject action is called, the object is successfully picked up. Since neither OpenVLA nor SpatialVLA was trained in the AI2-THOR environment, we relax this success threshold by setting the radius to 0.4 to account for discrepancies between training and evaluation environments. This loosened criterion ensures that small deviations in arm trajectories do not result in an immediate failure, thereby providing a fairer comparison of the models.
Figure 6 presents accumulated rewards on low-level manipulation tasks. Both models underperform relative to the Planner baseline, largely due to differences in robot arm configuration and the out-of-distribution nature of the visual inputs. The results show that SpatialVLA achieves slightly higher and more stable rewards than OpenVLA across long sequences.
Table 6 reports task success rates. Overall performance remains low, with both models struggling on manipulation tasks. We attribute these results primarily to two factors: (1) ManipulaTHOR’s egocentric view, which provides nearly identical images for fine-grained arm movements $(\Delta x,\Delta y,\Delta z)$ smaller than 0.01, making it difficult for VLA models to perceive subtle adjustments; and (2) discrepancies between the robot arm in ManipulaTHOR and those assumed by the VLA models, causing many predicted actions to fail in the physical simulation. We expect that fine-tuning on our dataset could provide an effective solution to bridge these gaps.
Appendix D Limitations
While $\infty$-Thor enables the generation of arbitrarily long trajectories, the diversity of environment layouts in AI2-THOR is inherently limited. This can lead to repetitive agent behaviors within certain scenes. For instance, compared to kitchen and living room scenes, bedroom scenes tend to involve fewer action types, mostly constrained to simple pick-and-place tasks within small spatial areas. Due to the limited scene size and low task diversity, we excluded bathroom scenes from our dev and test sets. In future work, we plan to integrate $\infty$-Thor with ProcTHOR (Deitke et al., 2022), which supports procedurally generated environments, enabling a broader range of dynamic and diverse scene configurations.
@@ Appendix D Limitations
Finally, inference with LLMs in long-context settings remains a computational bottleneck, particularly as context lengths approach 1M tokens. Since our long-horizon tasks require access to information spread throughout the entire trajectory, full-context inference becomes increasingly expensive, even with Context Parallelism. One promising direction is to equip agents with memory systems that selectively keep relevant information from prior steps, allowing the model to reason without reprocessing the full context at each step. Other architectural solutions, such as sparse attention mechanisms (Han et al., 2024; Jiang et al., 2024) or state-space models with linear-time inference (Gu & Dao, 2024), also hold potential for scalable long-context reasoning.